### v3

@@
[…]
Hagn et al.
[…]
Stochastic Team Formation and Routing
[…]
A Branch-Price-Cut-And-Switch Approach for Optimizing Team Formation and Routing for Airport Baggage Handling Tasks with Stochastic Travel Times
[…]
AdONE GRK 2201
[…]
AdONE GRK 2201
[…]
Department of Discrete Mathematics
[…]
In airport operations, optimally using dedicated personnel for baggage handling tasks plays a crucial role in the design of resource-efficient processes. Teams of workers with different qualifications must be formed, and loading or unloading tasks must be assigned to them. Each task has a time window within which it can be started and should be finished. Violating these temporal restrictions incurs severe financial penalties for the operator. In practice, various components of this process are subject to uncertainties. We consider the aforementioned problem under the assumption of stochastic travel times across the apron. We present two binary program formulations to model the problem at hand and propose a novel solution approach that we call Branch-Price-Cut-and-Switch, in which we dynamically switch between two master problem formulations. Furthermore, we use an exact separation method to identify violated rank-1 Chvátal-Gomory cuts and utilize an efficient branching rule relying on task finish times. We test the algorithm on instances generated based on real-world data from a major European hub airport with a planning horizon of up to two hours, 30 flights per hour, and three available task execution modes to choose from. Our results indicate that our algorithm is able to significantly outperform existing solution approaches. Moreover, an explicit consideration of stochastic travel times allows for solutions that utilize the available workforce more efficiently, while simultaneously guaranteeing a stable service level for the baggage handling operator.
[…]
@@ 1 Introduction
Baggage handling tasks, namely loading and unloading containers and bulk luggage, play an important role in aviation and airport operations. They are one of several ground handling tasks (Evler et al. 2021). Whenever a plane is arriving or scheduled for departure, a team of workers that operates several types of equipment needs to be formed and assigned to load or unload the aircraft. Said workers have different skill levels, i.e., qualifications for the types of equipment, such as tractors or high cargo loaders, they are allowed to operate. Depending on the airplane model and type, various equipment compositions can be used to execute the (un-)loading task, leading to different workforce requirements. For instance, large-sized airplanes typically have at least two cargo holes. These can be unloaded either sequentially or in parallel. For a more detailed description of the processes, see Dall’Olio and Kolisch (2023). In the following, whenever loading tasks are discussed, both loading and unloading processes are encompassed.
@@
Each type of equipment requires a certain skill level, where skill levels are ordered hierarchically. That is, a worker with skill level $k$ can execute any job requiring a skill level of $k$ or lower. Frequent delays in baggage claim processes and plane departures can result from improper team formation and task assignment decisions. This noticeably decreases passenger satisfaction and induces significant financial penalties for the baggage handling operator. Eurocontrol (2023) state that around 30% of all aircraft delays are caused by airline-related issues, to which baggage handling processes are attributed. Furthermore, depending on the airport and the considered planning horizon, baggage handling is subject to significant uncertainties regarding travel times of ground vehicles. Delays while traveling from one parking position to another, e.g., because of planes crossing the apron, are quite frequent and typically accumulate to a significant scope. In the following, we consider loading times to be deterministic, while travel times are assumed to be stochastic with known probability distributions. Moreover, to limit the potential financial penalties for the baggage handling operator, it is reasonable to demand that each task’s time window is satisfied with at least a predefined probability. Our work builds upon Dall’Olio and Kolisch (2023), which addresses the deterministic team formation and routing problem. The main contributions of this paper are:
@@
The remainder of this paper is structured as follows. Section 2 reviews the most relevant literature for our study. Section 3 provides a detailed problem description with the help of mathematical notation. In Section 4, we propose two binary programs that are used to model the problem at hand. In Section 5, we develop a Branch-Price-Cut-and-Switch solution approach and elaborate on its core components. Section 6 presents computational experiments aiming at assessing the proposed algorithm’s performance and the impact of stochasticity on optimal solutions. We summarize our findings in Section 7 and present several areas of future research.
2 Literature Review
The problem considered in this paper can be seen as a variant of the technician scheduling and routing problem (TSRP). TSRPs typically consist of a routing part that can be seen as a vehicle routing problem (VRP) and a scheduling or team formation part. Section 2.1 focuses on literature dealing with stochastic VRPs in a general context. Section 2.2 considers publications dealing with stochastic TSRP variants.
2.1 Literature on Stochastic Vehicle Routing
Recently, stochastic formulations of the vehicle routing problem with time windows (VRPTW) have seen a noticeable interest increase. In the following, we provide an overview of literature dealing with stochastic variants of the VRPTW. Oyola, Arntzen, and Woodruff (2018) summarize the most relevant literature on common types of stochastic VRPs. A taxonomy and overview of the space of VRPs can be found in Eksioglu, Vural, and Reisman (2009).
Generally, random variables in the stochastic VRPTW are assumed to be independent as stochastic dependency drastically complicates the calculation of distributions and their moments. While some publications focus on the VRPTW with stochastic demands (see Lee, Lee, and Park (2012), Zhang, Lam, and Chen (2016)), stochastic travel and service times are far more common. Stochasticity can be addressed in different ways. A common approach is to include chance constraints, which limit the probability of violating time windows. While Errico et al. (2018) connect chance constraints to the simultaneous satisfaction of all time windows, Ehmke, Campbell, and Urban (2015) as well as Li, Tian, and Leung (2010) limit the probability of violating the time window of each customer individually.
In vehicle routing problems, the distributions of start times at a given customer depend on the sequence of customers previously visited by the same vehicle and the distributions of travel and service times along the route. Furthermore, the computational effort required to calculate said distributions heavily depends on the type of time windows. Taş et al. (2014) assume that time windows are soft and travel times are independent, gamma distributed random variables. The authors propose an exact calculation of arrival time distributions by convolving finish time and travel time distributions and solve the problem using a Branch-Price-and-Cut approach. Because the convolution of gamma distributed random variables is gamma distributed, finish time distributions are again gamma distributed with known shape and scale parameter, which significantly simplifies their calculation.
If time windows are assumed to be hard, arrival time distributions need to be truncated at the start of each time window. Hence, distributional properties, such as moments or the existence of a closed-form expression, are not propagated along vehicle routes. Therefore, distributions can only be obtained by explicitly calculating the convolution of two distributions of arbitrary shapes, which is often not computationally tractable in reasonable time. Errico et al. (2018) assume service times to be stochastic, discretely distributed random variables and calculate start time distributions exactly. They use a Branch-Price-and-Cut algorithm to solve the problem and report computational results for four different variants of triangular distributions. Miranda and Conceição (2016) assume normally distributed service and travel times. They approximate service and travel time distributions by discretizing them on a finite set of values, whose count and values depend on the underlying distribution’s moments. This allows for an approximation of the finish time distribution of a given task by convolving the finish time distribution of the previous task with the aforementioned approximations of travel and service time distributions. They further improve their approach in Miranda, Branke, and Conceição (2018) using improved lower bounds for the approximation intervals of service and travel times. Ehmke, Campbell, and Urban (2015) show that the first and second moment of start time distributions can be calculated with little effort if both travel times and start times at the previous customer are normally distributed. While the former is usually not the case, the authors present computational results that indicate that their approach works well for normal, shifted gamma, and shifted exponential travel time distributions. Li, Tian, and Leung (2010) use stochastic simulation to derive estimates for start time distributions and solve the problem using tabu search.
Our considerations can be seen as a combination and extension of Errico et al. (2018) and Li, Tian, and Leung (2010), as we calculate start time distributions exactly but interpret chance constraints as a non-route-interdependent property and consider a stochastic objective function.
2.2 Literature on the stochastic TSRP
We first focus on literature on deterministic variants of the TSRP that share several key properties with the problem at hand. Pereira, Alves, and de Oliveira Moreira (2020) consider a multiperiod workforce scheduling and routing problem, where tasks do not have a fixed time window. However, precedence relationships need to be satisfied. The set of available teams and their characteristics, such as skill levels, are fixed and thus are not part of the decision space. The authors propose a mixed-integer model, which is shown to be computationally tractable only for small instances, and an ant-colony optimization heuristic. Çakırgil, Yücel, and Kuyzu (2020) examine a multi-objective workforce scheduling and routing problem, where teams consisting of workers with different skills need to be formed and task sequences need to be assigned to each team. Tasks can be started at any time, but need to be finished before their respective deadline. Workers can have different qualifications, but it is not possible for them to execute tasks that require a different qualification, i.e., downgrading is not possible. The authors developed a mixed-integer program and a 2-stage matheuristic to solve the problem. While the former proves to be inefficient for large instances, the latter is able to scale well with instance size. Li, Lim, and Rodrigues (2005) consider a manpower allocation problem with hard time windows. Similar to Çakırgil, Yücel, and Kuyzu (2020), workers with different qualifications need to be grouped into teams, which then execute a to-be-optimized sequence of tasks. Additionally, downgrading workers is not possible. The objective is to minimize a weighted sum of the total number of required workers and the total travel time. The authors develop a construction heuristic and a simulated annealing approach to solve the problem.
All of the publications previously discussed differ from our considerations in two aspects. First, they fix teams beforehand or make restrictive assumptions regarding team formation possibilities, such as the impossibility of downgrading or the existence of a single mode. Second, they focus on heuristic solution methods to solve medium- and large-sized instances. For an overview of further literature dealing with deterministic variants of the TSRP relevant to this study, we refer to Dall’Olio and Kolisch (2023).
While there is plenty of literature regarding the deterministic TSRP, very little research has been done on stochastic problem variants. In the following, we turn our focus on literature dealing with stochastic variants of the TSRP that are similar to ours.
Souyris et al. (2013) propose a robust formulation for the TSRP. Workers are assumed to be homogeneous and service times are stochastic. Each task must be started before a fixed deadline with a given probability. Moreover, each task must be executed by a single technician, thus team formation is not part of the model. The authors propose three different bounded uncertainty sets that limit the total service time delays per technician or client, respectively. The objective is to minimize the worst-case total delay and travel time. A branch-and-price approach is proposed and applied to real-world instances with 41 customers and 15 technicians.
Yuan, Liu, and Jiang (2015) address the TSRP in the context of health care workers and home health care services. Each task must be executed by a single worker. Furthermore, it is possible to downgrade workers to lower skill levels. Service times are assumed to be stochastic with a known probability distribution. The goal is to minimize the expected total travel costs, fixed costs of caregivers, service costs, and penalties for late arrival at customers. The authors use a branch-and-price approach to solve the problem exactly. Computational experiments assuming uniformly distributed service times and 25 to 50 customers indicate that the approach is able to provide very good results for small-sized instances. While the objective function and interpretation of stochasticity considered by Yuan, Liu, and Jiang (2015) is different from ours and several properties, such as the possibility of executing tasks using different modes and skill compositions, are missing, their solution approach exhibits similarities with the one proposed in the following and can be applied to the problem at hand. In Section 6.2, we compare both methodologies on a set of test instances.
Binart et al. (2016)
consider a TSRP with mandatory and optional tasks, where time windows are hard and fixed, and predefined teams must be used to execute tasks. Although not explicitly done, their modeling framework would allow considering multiple skill levels, downgrading, and multiple modes for single tasks. They assume that travel and service times are discretely distributed according to triangular distributions. Furthermore, the time windows of each task must be satisfied with a given probability. A 2-stage approach is proposed to heuristically solve the problem. First, a feasible skeleton solution is obtained using a generic MIP solver by optimally covering all mandatory tasks. Second, the first-stage solution is refined by inserting optional customers into the existing routes such that time window restrictions are not violated. Computational results on instances with 5 to 9 mandatory and 30 to 50 optional tasks indicate that the approach can yield good results in most cases. The problem considered by Binart et al. (2016) can be seen as a special case of the problem at hand where all technicians have the same skill level, each team formation represents one technician, and each task can be executed by each team formation. While the objective function and interpretation of stochasticity differ from ours, the models and solution approach proposed in the following can be applied to their problem.
Our work shows several similarities with Binart et al. (2016) and Yuan, Liu, and Jiang (2015), such as the stochasticity of travel times, the usage of exact solution methods, and the incorporation of several team properties such as multiple skill levels and modes. At the same time, our interpretation of service levels at individual tasks rather than entire solutions, the lack of assumptions on distributional properties, and the exact calculation of arrival times separate our considerations from previous approaches.
The following considerations are widely based on Dall’Olio and Kolisch (2023), which consider the problem of forming and routing worker teams for baggage handling tasks in a deterministic setting. We extend their deterministic problem description by assuming stochastic travel times. Furthermore, we advance their proposed solution algorithm by introducing cutting planes and additional branching strategies, improving the runtime efficiency of the pricing step, and describing an efficient way to handle infeasible solutions that is based on dynamically switching between two master problem formulations.
@@ 3 Problem Description
In the following section, we provide a detailed description of the problem considered in this publication with the help of mathematical notation, which will be used throughout the following sections. Table 1 provides an overview of general notation introduced in this section and used throughout the following sections.
Table 1: General notation
Sets and variables
$\mathcal{Q}$ Set of profiles
$\mathcal{Q}_{i}$ Set of profiles compatible with task $i$
$\mathrm{Q}^{\mathrm{D}}$ Set of disaggregated profiles $(q,s)$
$\mathcal{S}$ Set of skill compositions
$\mathcal{S}_{q}$ Set of skill compositions compatible with profile $q$
$\mathcal{I}$ Set of tasks
$\mathcal{I}^{r}$ Set of tasks executed by route $r$
$\mathcal{K}$ Set of worker skill levels
$\mathcal{T}$ Set of discrete time points
$B_{i,j}$ (finite) support of travel time delays for tasks $i$ and $j$ (or tasks and the depot)
$\Omega$ Set of all possible travel time realizations
$t$ random variable of stochastic travel times
$t_{i,j}$ random variable of stochastic travel times between nodes $i$ and $j$
Input Parameters
$N_{k}^{\mathrm{D}}$ Number of available workers with skill level $k$
$N_{k}$ Number of available workers with at least level $k$
$\xi_{q,k}$ Number of workers with at least level $k$ required by working profile $q$
$d$ Central worker depot
$\mathrm{ES}_{i}$ Earliest start time of task $i$
$\mathrm{LF}_{i}$ Latest finish time of task $i$
$\mathrm{LF}_{i}^{\mathrm{e}}$ Extended latest finish time of task $i$
$w_{i}$ Weight of task $i$
$p_{i,q}$ Processing time of task $i$ when undertaken with working profile $q$
$t^{\mathrm{det}}$ Vector of deterministic, best-case travel times
$\omega$ (single) realization of travel times
$\omega_{\mathrm{max}}$ worst-case realization of travel times
Model Parameters
$P_{i}(\cdot)$ Penalty function for time window violations at task $i$
$\mathbb{E}(c^{r})$ Expected cost of route $r$
$\mathrm{tl}^{r,q}$ Leave time from the depot for the team with profile $q$ executing route $r$
$\mathrm{tr}^{r,q}$ Return time to the depot for the team with profile $q$ executing route $r$
$S^{r,q}_{i}$ (Stochastic) start time of task $i$ in tour $r$ with profile $q$
$F^{r,q}_{i}$ (Stochastic) finish time of task $i$ in tour $r$ with profile $q$
$F^{r,q}_{i}(\omega_{\mathrm{max}})$ Worst-case finish time of task $i$ in tour $r$ with profile $q$
@@
We consider a set $\mathcal{I}=\{1,\dots,|\mathcal{I}|\}$ of tasks that have to be performed within a specified planning horizon. Each task corresponds to either a loading or unloading task of an incoming (or outgoing) flight. We note that, in the context of VRPs, tasks are typically referred to as customers or customer visits. A time window $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ is associated with each task $i\in\mathcal{I}$, describing the earliest and latest point at which service on an airplane can be started and should be finished, respectively.
@@
Initially, a workforce of prespecified size is located at a central depot denoted by $d$. In general, workers can be subdivided into skill levels $\mathcal{K}\coloneqq\{1,\dots,K\}$. Depending on their skill level, workers are only allowed to operate certain equipment. Skill levels are ordered hierarchically, i.e., a worker with skill level $k$ is able to operate any equipment requiring level $l\leq k$. Moreover, the amount of workers with skill level $\geq k$ that are available at the worker depot is fixed and defined as $N_{k}\coloneqq\underset{l=k}{\overset{K}{\sum}}N_{l}^{\mathrm{D}}$, where $N_{k}^{\mathrm{D}}\in\mathbb{N}_{>0}$ is the number of available workers with skill level $k$. We note that in practical instances, higher-skilled workers tend to be rather scarce and hard to obtain as they require lengthy and expensive training. A route is a sequence $r\coloneqq\left(d,v_{1},\dots,v_{l},d\right)$ with $l\geq 1$ and $v_{i}\in\mathcal{I}$ for all $i=1,\dots,l$. Each team route $(r,q)$ has time instants $\mathrm{tl}^{r,q}$ and $\mathrm{tr}^{r,q}$ at which the team leaves the depot and returns to it, respectively. The set of tasks executed by route $r$ is denoted by $\mathcal{I}^{r}\coloneqq\{v_{1},\dots,v_{l}\}$. Furthermore, each route $r$ is assigned a profile $q=\left(\xi_{q,k}\right)_{k\in\mathcal{K}}\in\mathcal{Q}$ which is used to execute the route, where $\xi_{q,k}\in\mathbb{N}_{\geq 0}$ denotes the number of required workers of skill level greater than or equal to $k$ and $\mathcal{Q}$ is the set of available profiles. Because each type of aircraft requires different prerequisites and equipment for baggage handling tasks, only a subset $\mathcal{Q}_{i}\subseteq\mathcal{Q}$ of profiles can be used to execute a given task $i\in\mathcal{I}$. Depending on the profile $q\in\mathcal{Q}_{i}$, executing task $i$ takes $p_{i,q}\in\mathbb{N}_{>0}$ units of time.
@@
By construction, a profile $q\in\mathcal{Q}$ is well-defined by its aggregated workforce requirements, i.e., lower bounds for the skill levels of required workers, but does not contain any information about their actual skill levels. To include this information, we define a skill composition as a vector $s\in\mathbb{N}^{|\mathcal{K}|}$ and say that a skill composition $s=(s_{q,k})_{k\in\mathcal{K}}$ be can assigned to a profile $q\in\mathcal{Q}$ if
@@
as the set of disaggregated profiles, i.e., profiles enhanced by information on the skill levels of workers used to assemble a team with profile $q$. Given a route $r$, we define the set of profiles that can be used to execute $r$ as $\mathcal{Q}^{r}\coloneqq\underset{i\in\mathcal{I}^{r}}{\bigcap}\mathcal{Q}_{i}$. An example illustrating the relationship between profiles and skill compositions is presented in Section 4.2.
@@
After finishing a task $i$, a team with profile $q$ can either return to the depot $d$ and become available for regrouping or directly continue with another task $j$. Traveling between locations $i,j\in\mathcal{I}\cup\left\{d\right\}$, where $i$ and $j$ are either two tasks or one task and the depot, takes at least $t_{i,j}^{\mathrm{det}}\in\mathbb{N}_{\geq 0}$ units of time and does not depend on the profile. We assume travel times to be symmetric and time-independent. In practice, these travel times are subject to delays caused by unexpected, exogenous factors, such as aircrafts crossing the apron or local traffic congestion. In such situations, a common approach is to solve a multi-stage stochastic problem, where an initial solution is constructed and, once more information on stochastic parameters such as travel times becomes available over time, is readjusted if needed. However, as ground vehicle traffic is rather unpredictable and an aircraft’s movements are typically only determined moments before the plane starts moving, it is not possible to avoid the usage of congested parts of the apron by re-adjusting the previously derived solution, i.e., by changing the sequences of tasks that should be executed by the same team or assigning a new team to a given task. To formalize this, we define non-negative, finite supports for travel time delays for each pair $(i,j)$ of tasks (or tasks and the depot) by
$B_{i,j}\subset\mathbb{N}_{\geq 0},\ |B_{i,j}|<\infty\quad\forall(i,j)\in E$

### v4

@@
Abstract
In airport operations, optimally using dedicated personnel for baggage handling tasks plays a crucial role in the design of resource-efficient processes. Teams of workers with different qualifications must be formed, and loading or unloading tasks must be assigned to them. Each task has a time window within which it can be started and should be finished. Violating these temporal restrictions incurs severe financial penalties for the operator. In practice, various components of this process are subject to uncertainties. We consider the aforementioned problem under the assumption of time-dependent stochastic travel times across the apron. We present two binary program formulations to model the problem at hand and propose a novel solution approach that we call Branch-Price-Cut-and-Switch, in which we dynamically switch between two master problem formulations. Furthermore, we use an exact separation method to identify violated rank-1 Chvátal-Gomory cuts and utilize an efficient branching rule relying on task finish times. We test the algorithm on instances generated based on real-world data from a major European hub airport with a planning horizon of up to two hours, 30 flights per hour, and three available task execution modes to choose from. Our results indicate that our algorithm is able to significantly outperform existing solution approaches. Moreover, an explicit consideration of stochastic travel times allows for solutions that utilize the available workforce more efficiently, while simultaneously guaranteeing a stable service level for the baggage handling operator.
Highlights
  • 1.
    Extension of the problem of forming and routing worker teams for baggage handling tasks to the setting of time-dependent stochastic travel times
  • 2.
    Proposal of a novel branch-price-cut-and-switch scheme, which dynamically utilizes and switches between two master problem formulations depending on the problem characteristics
  • 3.
    Computational tests show a significant outperformance of comparable exact approaches on a test instance set based on real-world data
  • 4.
    Explicit consideration of stochastic travel times ensures minimum service levels and improved safety buffers compared to the usage of deterministic estimators for travel times, as well as
keywords:
@@ 1 Introduction
Baggage handling tasks, namely loading and unloading containers and bulk luggage, play an important role in aviation and airport operations. They are one of several ground handling tasks (evler2021airline). Baggage handling is usually performed by dedicated personnel using dedicated equipment. Furthermore, there are no dependencies between baggage handling and other ground handling tasks, making it possible to consider baggage (un)loading in isolation from all other ground handling tasks. Typically, a central planner needs to assemble worker teams and decide which aircraft they (un)load and in which sequence they do so.
Workers have different skill levels, i.e., qualifications for the types of equipment, such as tractors or high cargo loaders, they are allowed to operate. Depending on the airplane model and type, various equipment compositions can be used to execute the (un-)loading task, leading to different workforce requirements. For instance, large-sized airplanes typically have at least two cargo holes. These can be unloaded either sequentially or in parallel. Each aircraft can be (un)loaded in several such modes, where the use of additional equipment implies the requirement of additional workers and typically leads to shorter execution times. Such equipment compositions are also called modes or (team) formations and are derived from standardized (un)loading procedures, which are defined in the respective aircraft manual. Each flight is assigned a time window, within which the baggage handling needs to take place. A central planner can reduce the execution time of a task by selecting a mode that either requires more workers in general or that requires higher-skilled workers. Because workforce is a scarce resource in baggage handling, the selection of the fastest possible mode for each incoming and outgoing flight is typically not feasible. At the same time, selecting slow modes too frequently can lead to the violation of time windows, causing departure delays and delayed baggage pick-up for arriving flights. Because such disruptive delays can significantly reduce customer satisfaction and incur financial penalties for the baggage handling operator, proper team formation and sequencing decisions by the central planner are crucial. Thus, a well-informed and efficient assignment of workers to flights and sequencing of flights can greatly benefit customer satisfaction and reduce fines paid by the operator. In the following, whenever loading tasks are discussed, both loading and unloading processes are encompassed.
@@
coda2023q2 state that around 30% of all aircraft delays are caused by airline-related issues, to which baggage handling processes are attributed. Furthermore, depending on the airport and the considered planning horizon, baggage handling is subject to significant uncertainties regarding travel times of ground vehicles. Delays while traveling from one parking position to another, e.g., because of planes crossing the apron, are quite frequent and typically accumulate to a significant extent. In the following, we consider loading times to be deterministic, while travel times are assumed to be stochastic with known probability distributions. Moreover, to limit the potential financial penalties for the baggage handling operator, it is reasonable to demand that each task’s time window is satisfied with at least a predefined probability.
If the decision of how to form teams and how to assign them to incoming or outgoing flights is done improperly, one might end up with a shortage of workers or unnecessarily long execution times. This in turn causes delays within the baggage handling process, which frequently propagate for several hours and lead to increased waiting times at the baggage claim area or departure delays. Both of these inefficiencies significantly impact passenger satisfaction. Optimizing the planning and processing of baggage (un)loading tasks can help to avoid excessively large delays or even eliminate them completely. For this reason, we consider the problem of forming and routing worker teams for baggage handling tasks under stochastic travel times, such that each task is executed by exactly one team and the available workforce is not exceeded. The objective function minimizes the sum of weighted expected finish times of all tasks and a penalty term for exceeding time windows. While the former aims to introduce buffers that can be used to compensate unforeseen delays during baggage handling, the latter considers financial fines paid by the baggage handling operator in case of delays, as well as an increasing chance of passengers missing connecting flights.
Our work builds upon dallolio_temp, which addresses the deterministic team formation and routing problem. The main contributions of this paper are:
@@
The remainder of this paper is structured as follows. Section 2 provides a detailed problem description with the help of mathematical notation. Section 3 reviews the most relevant literature for our study. In Section 4, we propose two binary programs that are used to model the problem at hand. In Section 5, we develop a Branch-Price-Cut-and-Switch solution approach and elaborate on its core components. Section 6 presents computational experiments aiming at assessing the proposed algorithm’s performance and the impact of stochasticity on optimal solutions. We summarize our findings in Section 7 and present several areas of future research.
@@ 2 Problem Description
In the following section, we provide a detailed description of the problem considered in this publication with the help of mathematical notation, which will be used throughout the following sections. An overview over all used notation can be found in A.
@@
We consider a set $\mathcal{I}=\{1,\dots,|\mathcal{I}|\}$ of tasks that have to be performed within a specified planning horizon. The planning horizon is discretized into a set $\mathcal{T}$ of equidistant time steps. Each task corresponds to either a loading or unloading task of an incoming (or outgoing) flight. We note that, in the context of VRPs, tasks are typically referred to as customers or customer visits. A time window $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ is associated with each task $i\in\mathcal{I}$, describing the earliest and latest point at which service on an airplane can be started and should be finished, respectively.
@@
Initially, a workforce of prespecified size is located at a central depot denoted by $d$. In general, workers can be subdivided into skill levels $\mathcal{K}\coloneqq\{1,\dots,K\}$. Depending on their skill level, workers are only allowed to operate certain equipment. For instance, aircraft typically carry bulk luggage, containers, and/or pallets. While skill level $1$ only encompasses basic apron clearance, skill level $2$ allows for the operation of loader-transporters to unload containers. Heavy equipment used to unload multiple containers or pallets at once, such as high-loaders, requires a qualification of level $3$. Skill levels are ordered hierarchically, i.e., a worker with skill level $k\in\mathcal{K}$ is able to operate any equipment requiring level $l$ not greater than $k$. Moreover, the amount of workers with skill level greater than or equal to $k$ that are available at the worker depot is fixed and defined as $N_{k}\coloneqq\underset{l=k}{\overset{K}{\sum}}N_{l}^{\mathrm{D}}$, where $N_{k}^{\mathrm{D}}\in\mathbb{N}_{>0}$ is the number of available workers with skill level $k\in\mathcal{K}$. We note that in practical instances, higher-skilled workers tend to be rather scarce and hard to obtain as they require lengthy and expensive training. A route is a sequence $r\coloneqq\left(d,v_{1},\dots,v_{l},d\right)$ with $l\geq 1$ and $v_{i}\in\mathcal{I}$ for all $i=1,\dots,l$. The set of tasks executed by route $r$ is denoted by $\mathcal{I}^{r}\coloneqq\{v_{1},\dots,v_{l}\}$. Furthermore, each route $r$ is assigned a profile $q=\left(\xi_{q,k}\right)_{k\in\mathcal{K}}\in\mathcal{Q}$ which is used to execute the route, where $\xi_{q,k}\in\mathbb{N}_{\geq 0}$ denotes the number of required workers of skill level greater than or equal to $k$ and $\mathcal{Q}$ is the set of available profiles. Because each type of aircraft requires different prerequisites and equipment for baggage handling tasks, only a subset $\mathcal{Q}_{i}\subseteq\mathcal{Q}$ of profiles can be used to execute a given task $i\in\mathcal{I}$. For example, one can choose to either unload multiple cargo holes sequentially or in parallel, yielding different workforce requirements and thus different profiles. For larger aircraft, one can also choose between the usage of loader-transporters or high-loaders, leading to additional suitable profiles. Depending on the profile $q\in\mathcal{Q}_{i}$, executing task $i$ takes $p_{i,q}\in\mathbb{N}_{>0}$ units of time. The execution time of a task depends on several external factors, such as the number of containers, pallets, and the amount of bulk luggage that are to be handled. Each team route $(r,q)\in\mathcal{R}$, where $\mathcal{R}$ is the set of tuples of routes and suitable profiles, has time steps $\mathrm{tl}^{r,q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{T}}$ and $\mathrm{tr}^{r,q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{T}}$ at which the team leaves the depot and returns to it, respectively.
@@
By construction, a profile $q\in\mathcal{Q}$ is well-defined by its aggregated workforce requirements, i.e., lower bounds for the skill levels of required workers, but does not contain any information about their actual skill levels. To include this information, we define a skill composition as a vector $s\in\mathbb{N}^{|\mathcal{K}|}$ and say that a skill composition $s=(s_{q,k})_{k\in\mathcal{K}}$ can be assigned to a profile $q\in\mathcal{Q}$ if
@@
as the set of disaggregated profiles, i.e., profiles enhanced by information on the skill levels of workers used to assemble a team with profile $q$. Given a route $r$, we define the set of profiles that can be used to execute $r$ as $\mathcal{Q}^{r}\coloneqq\underset{i\in\mathcal{I}^{r}}{\bigcap}\mathcal{Q}_{i}$.
Example 1.
$\xi_{q}\coloneqq\left(\xi_{q,k})\right)_{i=1,2,3}=\left(\begin{array}[]{c}3\\ 2\\ 1\\ \end{array}\right).$
Furthermore, we assume that we have an unlimited workforce available for each skill level. In practice, a team of profile $q$ could operate as follows. First, one worker operates a high-loader, which unloads and moves containers onto dollies. This requires skill level $3$, as the handling of high-loaders requires lengthy prior training. Second, one worker moves a conveyor belt into position and unloads bulk luggage into dollies. Because conveyor belts are significantly easier to handle than high-loaders, a skill level of at least $2$ is required for this job. Third, one worker unloads bulk luggage from the aircraft onto the conveyor belt. This is a rather simple task and can be executed by any worker of skill level $1$ or higher.
Table 1: Set of possible skill compositions for profile $q$
$s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s^{i}_{q,1}$ 1 0 0 1 0 $s^{i}_{q,2}$ 1 2 1 0 0 $s^{i}_{q,3}$ 1 1 2 2 3
Table 1 contains the set $\mathcal{S}_{q}=\left\{s_{1},s_{2},s_{3},s_{4},s_{5}\right\}$ of skill compositions that can be used to assemble a team with profile $q$. We note that each skill composition $s_{i}$ is uniquely defined by its worker usage $\left(s^{i}_{q,k}\right)_{k=1,2,3}$ per skill level. For instance, the third column in Table 1 describes skill composition $s_{2}$, which uses two workers of skill level 2 and one worker of skill level 3 to assemble a team with profile $q$. Note that this implies that one worker with level $2$ is assigned the job of unloading bulk luggage from the aircraft onto the conveyor belt, i.e., he is downgraded to a job below his actual skill level. Furthermore, a team with profile $q$ always occupies the same workforce on an aggregated level, i.e., such a team always requires 3 workers of at least level 1, 2 workers of at least level 2, and one worker of level 3. However, depending on the selected skill composition, the workforce occupation per individual level differs. For instance, while $s_{4}$ occupies two workers of level 3, $s_{5}$ occupies only one worker of level 3. From a modelling point of view, considering workforce requirements on an aggregated level is often sufficient to create a feasible plan. However, on some occasions, this aggregated consideration leads to an infeasible plan, which can be handled by explicitly specifying and considering skill compositions. In total, team route $(r,q)$ can be used to derive five unique team routes $\left\{(r,q,s_{i}):\ i=,1\dots,5\right\}$ in $\mathcal{R}^{\mathrm{D}}$. Thus, the number of skill compositions typically significantly exceeds the number of profiles.
@@
After finishing a task $i$, a team with profile $q$ can either return to the depot $d$ and become available for regrouping or directly continue with another task $j$. Traveling between locations $i,j\in\mathcal{I}\cup\left\{d\right\}$, where $i$ and $j$ are either two tasks or one task and the depot, takes at least $t_{i,j}^{\mathrm{det}}\in\mathbb{N}_{\geq 0}$ units of time and does not depend on the profile. We assume that travel times are stochastic and time-dependent. For this purpose, we decompose the planning horizon into subsequent time bins of equal length. Formally, we define a disjoint decomposition $\mathcal{B}\coloneqq\left\{B_{1},\dots,B_{m}\right\}$ with $m\in\mathbb{N}$ into sets $B_{i}\subset\mathcal{T}$ of equal cardinality and length. For instance, given a 2 hour planning horizon $\mathcal{T}=\{1,\dots,120\}$ and a time bin length of 15 minutes, a decomposition is then given by $\mathcal{B}_{i}=\{t_{15i+1},\dots,t_{15i+15}\}$ for $i=0,\dots,7$. Given a time bin $\mathcal{B}_{i}\in\mathcal{B}$, we assume that travel times are independent. Note that travel times across different time bins are, in general, dependent. Practically speaking, a time bin encompasses a period of stable traffic across the apron, in which uncertainties on travel times predominantly arise from unsystematic causes that affect individual pathways. Such independent delays can, for instance, be caused by aircraft crossing the apron or local traffic congestion. In such situations, a common approach is to solve a multi-stage stochastic problem, where an initial solution is constructed and, once more information on stochastic parameters such as travel times becomes available over time, is readjusted if needed. However, as ground vehicle traffic is rather unpredictable and an aircraft’s movements are typically only determined moments before the plane starts moving, it is not possible to avoid the usage of congested parts of the apron by re-adjusting the previously derived solution, i.e., by changing the sequences of tasks that should be executed by the same team or assigning a new team to a given task. To formalize this, we define non-negative, finite supports for travel time delays for each time bin $\mathcal{B}_{k}\in\mathcal{B}$ and each pair $(i,j)\in E$ of tasks (or tasks and the depot) by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{i,j}^{k}}\subset\mathbb{N}_{\geq 0},\ |{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{i,j}^{k}}|<\infty$